could you please explain why are the answers to both sides of this equation equal? When N = 1500
100 + 0.05N = 70 + 0.07N

Answer by Alan3354(69443)

could you please explain why are the answers to both sides of this equation equal? When N = 1500
100 + 0.05N = 70 + 0.07N
Sub 1500 for N
------------
100 + 0.05*1500 =? 70 + 0.07*1500
100 + 75 =? 70 + 105
175 = 175
They are equal at N = 1500.